22jan16: alfa bm6b instabilities
Jason hessels (p3055)
reported problems with alfa bm 6. he reported problems with data
- 08dec15,17dec15, and 26dec15.
I finally got around to looking into this using the 20jan16
Processing the data:
- 1 millisecond sampling, winking cal (20ms on, 20ms off)
- 128 channel spectra, full stokes (after decimating in freq).
- i used the 172 mhz channel centered at 1450 Mhz.
- driving za from 2 to 19.5 degrees za (basket weaving).
- a nights run lasted for about 1 hour.
- i read in the data, averaged the 172Mhz 1450cfr band giving 1
total power sample per millisecond
- I then average over the 40millisec cal cycle (throwing out the
transition points) giving 1 total power sample each 40
The first set of plots shows the
total power vs time for bm0,1, and beam 6 (.ps) (.pdf)
- black in pola, red is polb
- the data sampling ends up being 40 millisecs/point
- top frame: beam 0
- middle frame: beam1
- bottom frame: beam6
- you can see the system temperature change with za
- beam 6 has stability problems when it is at low zenith
The 2nd set of plots, shows beam 6
vs za and time (.ps) (.pdf)
- black is polA, red is polB
- Top frame: total power vs za for then 1 hour run (2.5 times
up/down in za)
- polB is going unstable below about za=6
- middle frame: blowup show the za vs time for the first
pass through za=2
- bottom frame: this is the corresponding total power vs time
for the za track in the middle frame.
- the dashed green lines enclose when the za was below 3 deg.
- coming down the instabilities started around 3deg za. On the
way up they continues until around 4.5 deg za.
The za dependence looks like there is something loose.
- The receivers are level at za=10, tilted -10deg at
za=0,and tilted +10deg at za=20deg.
On 22jan16 we looked at the bias voltages and saw that beam 6a
stage 1 biases were jumping around
- This was happening at za's above the 4.2 degrees of the plots
We monitor the alfa bias's (usually once every 5 minutes). But..
during pulsar experiments, the bias monitoring has to be turned
off so it doesn't affect the observations (reading the bias's
causes glitches in the bias's output).
The plots below show the measured
Drain Current Id for stage 1of the alfa amplifiers for
01jan-22jan16 (.ps) (.pdf)
- Top frame: PolA stage 1 Id drain currents for amps.
- Each beam is color coded. beam 6 is light blue
- The dropouts are when the monitoring was turned off
- Bottom frame: polB stage 1 Id drain current
- b, 6b Id jumped up to 17 milliamps on 17jan and 20jan16 (and
stayed there till end of plot.. 22jan16).
- The 20jan16 galfacts data was taken while Id was at 17
I went back and looked at the alfa bias monitoring during jason's
runs on 08,17,26 dec15. Unfortunately the bias monitoring was
turned off during this time period (the operators turn off the
monitoring, but it doesn't always get turned back on when alfa
observing is done).
- The instabilities in beam 6 is coming from 6b. PolA is ok
- The problems are coming from the bias values jumping around
- The original za dependence from 20jan16 data points to
something loose (because of the changing tilt with za)
- but on 22jan16 it looked like it was happening on
other za's outside the 0-4.2za range.
- We need to find out if the problem is inside or outside the